OpenStudy (anonymous):

A satellite is projected vertically upwards from an earth station. At what height above the earth's surface will the force on the satellite due to the earth be reduced to half its value at the earth station(radius of earth 6400km)

OpenStudy (anonymous):

Whats wrong with my approach? \[GMm/2r" ^{2}=GMm/r ^{2}\] \[r=\sqrt{2(4096*10^{10})}\] r=8960000 Distance from earth's surface=8960000-6400000m=2560km The ans is supposed to be 2650km

OpenStudy (mani_jha):

Is r the earth's radius?

OpenStudy (anonymous):

distance of separation of centre of earth and satellite when force is half its initial value

OpenStudy (mani_jha):

If r is the new distance, and R is the radius of the earth(i.e the initial separation) then: \[r =\sqrt{2R}\] Why have you substituted R=4096X10^10? R=6400km

OpenStudy (anonymous):

GMm cancel so sqrt(2r"^2)=r ?

OpenStudy (anonymous):

F(R+h)=0.5F(R) 1/(R+h)^2= 0.5/R^2 2R^2=(R+h)^2

OpenStudy (anonymous):

where exactly did i go wrong?

OpenStudy (anonymous):

plz do it with my method and tell me what i did wrong

OpenStudy (anonymous):

even with this method i get the same ans

OpenStudy (anonymous):

give some time... you didnt,i think so!!! 2R^2=(R+h)^2 81920000=40960000+12800h+h2 h2+12800h+40960000=0 check this!!!!

OpenStudy (anonymous):

\[h=(\sqrt{2}-1)r\] h=(.4)(6400)=2560

OpenStudy (anonymous):

is this the solution to my quadratic ????

OpenStudy (anonymous):

i think the answer is wrong,maybe!!!

OpenStudy (anonymous):

this is from hc verma i dont think hes ever wrong

OpenStudy (anonymous):

no there are some errors in this book also, i too practised from his book sometimes...

OpenStudy (mani_jha):

Yes that is correct, shayan. The answer could be quadratic. You can google H.C. Verma solutions to check this. Use a calculator, Sarkar

OpenStudy (anonymous):

mate which calc...wolf has a different solution

OpenStudy (anonymous):

Quoting from HC Verma sol h=0.414*6400=2649.6

OpenStudy (anonymous):

but what exactly is wrong with the first method?

OpenStudy (anonymous):

please give link to the solutions...

OpenStudy (mani_jha):

Must be a calculation mistake.

OpenStudy (anonymous):

google search download cant view

OpenStudy (anonymous):

i agree @Mani_Jha

OpenStudy (anonymous):

the solutions books takes a more accurate value of sqrt2 but what about my method what went wrong?

OpenStudy (anonymous):

oh so i guess im supposed to take the more accurate value of root 2 as well

OpenStudy (mos1635):

Whats wrong with my approach? GMm/2r"2=GMm/r2 r=2(4096∗1010)−−−−−−−−−−−√ i think i got it!! what is 4096*10^10 ????

OpenStudy (mos1635):

it should be 6400exp3

OpenStudy (anonymous):

where did root go?

OpenStudy (mos1635):

r=R*sqr2 r=6400exp3* 1.4142

OpenStudy (anonymous):

i got the ans by taking 1.414 for root2 but who in the holy world does that?

OpenStudy (mos1635):

r=9050966m R+h=9050966m h=2650966m

OpenStudy (anonymous):

everyone does that...

OpenStudy (anonymous):

lol i never do

OpenStudy (mani_jha):

r=sqrt(2(4096∗1010)−−−−−−−−−−−√ r=8960000 Check this with calculator

OpenStudy (anonymous):

then you wont get an accurate answer,lol

OpenStudy (anonymous):

no, i wont @Mani_Jha

OpenStudy (anonymous):

thx people

OpenStudy (mani_jha):

There's a calculation mistake there.

OpenStudy (anonymous):

@Mani_Jha -what are you typing,man -type fast....

OpenStudy (mani_jha):

Why? Got to catch a flight? :P

OpenStudy (anonymous):

rather, gotta catch a bus!!!!

OpenStudy (mani_jha):

Have a safe journey! Did u find your mistake, shayan?

OpenStudy (anonymous):

yeah i did i think i mentioned that